# S1-NexusAgent: a Self-Evolving Agent Framework for Multidisciplinary Scientific Research

### v1

@@ Abstract
Modern scientific research relies on large-scale data, complex workflows, and specialized tools, which existing LLMs and tool-based agents struggle to handle due to limitations in long-horizon planning, robust goal maintenance, and continual learning from execution. To address these issues, in this work, we propose S1-NexusAgent, a self-evolving agent framework designed for multidisciplinary scientific research. S1-NexusAgent adopts a hierarchical Plan-and-CodeAct execution paradigm, decoupling global scientific planning from subtask-level tool execution through a dual-loop architecture, thereby enabling stable modeling of complex research workflows. The system natively supports the Model Context Protocol (MCP), integrates up to thousands of cross-disciplinary scientific tools, and achieves efficient orchestration of heterogeneous research tools via intention-aware dynamic tool retrieval and hot-plug mechanisms. To address long-context and large-scale data challenges in scientific settings, S1-NexusAgent introduces object-reference-based sparse context management, which enables sub-task context isolation and intermediate result compression. Building on this, a Critic Agent automatically evaluates complete execution trajectories and distills high-quality research paths into reusable Scientific Skills, forming a closed loop for continuous self-evolution, which is valuable for sustainable and long-horizon scientific research. Experiments on authoritative scientific benchmarks involving long-horizon planning and complex specialized tool orchestration, including biomini-eval (biology), ChemBench (chemistry), and MatSciBench (material science), demonstrate that S1-NexusAgent achieves state-of-the-art performance, validating its effectiveness and generalization capability in complex scientific tasks.
@@ 1.1 Domain-Specific Scientific Agents
A large body of recent work builds domain-specialized scientific agents that encode substantial task priors for a particular field, such as molecular design and optimization, materials discovery/characterization, and biomedical discovery workflows. Representative systems include chemistry and materials agents like ChemCrow [Bran et al., 2024], CACTUS [McNaughton et al., 2024], IR-Agent [Noh et al., 2025], HoneyComb [Zhang et al., 2024], MatPilot [Ni et al., 2024], MAPPS [Zhou et al., 2025], and others, as well as life-science agents spanning drug discovery, genomics/bioinformatics, and protein/CRISPR applications (e.g., DrugAgent [Inoue et al., 2024]/DrugPilot [Li et al., 2025a], Biomni [Huang et al., 2025], BioDiscoveryAgent [Roohani et al., 2024], CellVoyager [Alber et al., 2025], GeneGPT [Jin et al., 2024], CRISPR-GPT [Huang et al., 2024], etc.). These agents often achieve strong performance by aligning planning and tool use with a narrowly defined workflow or domain toolchain.
@@ 1.3 Code-enhanced and programmatic scientific agents
Another line of related work emphasizes code as a first-class action and planning medium for scientific agents. Rather than relying solely on natural-language reasoning, these systems generate executable artifacts—such as programs, scripts, or structured workflows—to drive scientific computation, data analysis, and experimental validation, improving precision, reproducibility, and controllability. Representative systems include AlphaEvolve [Novikov et al., 2025], which performs iterative program synthesis and evaluation for scientific optimization; workflow-oriented agents such as AIGS [Liu et al., 2024] and ORGANA [Darvish et al., 2025], which synthesize executable pipelines for scientific processes; and tool-augmented analysis agents such as Chemist-X [Chen et al., ] and Biomni [Huang et al., 2025], which iteratively generate and execute code to validate hypotheses and refine results. In these systems, planning is often encoded implicitly through program structure or explicitly via intermediate program representations, with execution serving as the primary grounding mechanism.
@@ 6 Discussion
To rigorously evaluate the individual contributions of our methodology, we conduct systematic ablation studies along two complementary dimensions, with all experiments evaluated on the Biomni Eval-1 benchmark: (1) the effectiveness of the proposed Inner–Outer Dual-Loop architecture, and (2) the impact of staged training strategies, comparing supervised fine-tuning (SFT) and reinforcement learning (RL), on agentic execution quality.
For the Dual-Loop architecture ablation, all architectural variants are evaluated on Biomni-Eval-1 using Qwen3-8B-SciCodeAct—a model specifically fine-tuned for scientific programming—as a fixed backbone, ensuring that performance differences are solely attributable to architectural design.
For the training strategy ablation, we likewise conduct evaluation on Biomni-Eval-1, fixing the backbone to the base Qwen3-8B model and comparing Only SFT against SFT + RL under identical conditions, thereby isolating the effect of reinforcement learning on agentic execution.
@@ 6.1 Ablation on the Dual-Loop Architecture
Table 1: Ablation study of the Dual-Loop Architecture on Biomni-Eval-1. Results compare isolated loop configurations against the integrated framework. Best results are highlighted in bold.
@@ 6.2 Ablation on RL-based Staged Training
We further analyze how staged learning enhances the execution quality of the CodeAct module. In this context, staged training refers to an SFT initialization that enforces protocol adherence, followed by RL-based policy optimization that refines decision-making under uncertainty. Using the same Qwen3-8B backbone, we compare Only SFT against SFT + RL.
@@
Table 2: Ablation study of RL-based Staged Training on Biomni-Eval-1. All configurations utilize the Qwen3-8B-SciCodeAct backbone. Best results are highlighted in bold.
@@ 7 Conclusion
Extensive experiments on Biomni-Eval, ChemBench, and MatSciBench show that S1-NexusAgent achieves state-of-the-art performance across life sciences, chemistry, and materials science, validating both its effectiveness and generalization capability in complex scientific tasks. Case studies further illustrate how the dual-loop architecture and code-centric execution paradigm enable interpretable, adaptive, and resilient scientific workflows in realistic research scenarios.

### v2

@@ Abstract
Modern scientific research relies on large-scale data, complex workflows, and specialized tools, which existing LLMs and tool-based agents struggle to handle due to limitations in long-horizon planning, robust goal maintenance, and continual learning from execution. To address these issues, in this work, we propose S1-NexusAgent, a self-evolving agent framework designed for multidisciplinary scientific research. S1-NexusAgent adopts a hierarchical Plan-and-CodeAct execution paradigm, decoupling global scientific planning from subtask-level tool execution through a dual-loop architecture, thereby enabling stable modeling of complex research workflows. The system natively supports the Model Context Protocol (MCP), integrates up to thousands of cross-disciplinary scientific tools, and achieves efficient orchestration of heterogeneous research tools via intention-aware dynamic tool retrieval and hot-plug mechanisms. To address long-context and large-scale data challenges in scientific settings, S1-NexusAgent introduces object-reference-based sparse context management, which enables sub-task context isolation and intermediate result compression. Building on this, a Critic Agent automatically evaluates complete execution trajectories and distills high-quality research paths into reusable Scientific Skills, forming a closed loop for continuous self-evolution, which is valuable for sustainable and long-horizon scientific research. Experiments on authoritative scientific benchmarks involving long-horizon planning and complex specialized tool orchestration, including biomini-eval (biology), ChemBench (chemistry), and MatSciBench (material science), demonstrate that S1-NexusAgent achieves state-of-the-art performance, validating its effectiveness and generalization capability in complex scientific tasks. The code of S1-NexusAgent is available at https://github.com/CASIA-LM/S1-NexusAgent.
@@ 1.1 Domain-Specific Scientific Agents
A large body of recent work builds domain-specialized scientific agents that encode substantial task priors for a particular field, such as molecular design and optimization, materials discovery/characterization, and biomedical discovery workflows [Ren et al., 2026]. Representative systems include chemistry and materials agents like ChemCrow [Bran et al., 2024], CACTUS [McNaughton et al., 2024], IR-Agent [Noh et al., 2025], HoneyComb [Zhang et al., 2024], MatPilot [Ni et al., 2024], MAPPS [Zhou et al., 2025], and others, as well as life-science agents spanning drug discovery, genomics/bioinformatics, and protein/CRISPR applications (e.g., DrugAgent [Inoue et al., 2024]/DrugPilot [Li et al., 2025a], Biomni [Huang et al., 2025], BioDiscoveryAgent [Roohani et al., 2024], CellVoyager [Alber et al., 2025], GeneGPT [Jin et al., 2024], CRISPR-GPT [Huang et al., 2024], etc.). These agents often achieve strong performance by aligning planning and tool use with a narrowly defined workflow or domain toolchain.
@@ 1.3 Code-enhanced and programmatic scientific agents
Another line of related work emphasizes code as a first-class action and planning medium for scientific agents. Rather than relying solely on natural-language reasoning, these systems generate executable artifacts—such as programs, scripts, or structured workflows—to drive scientific computation, data analysis, and experimental validation, improving precision, reproducibility, and controllability. Representative systems include AlphaEvolve [Novikov et al., 2025], which performs iterative program synthesis and evaluation for scientific optimization; workflow-oriented agents such as AIGS [Liu et al., 2024] and ORGANA [Darvish et al., 2025], which synthesize executable pipelines for scientific processes; and tool-augmented analysis agents such as Chemist-X [Chen et al., 2023] and Biomni [Huang et al., 2025], which iteratively generate and execute code to validate hypotheses and refine results. In these systems, planning is often encoded implicitly through program structure or explicitly via intermediate program representations, with execution serving as the primary grounding mechanism.
@@ 6 Discussion
To rigorously evaluate the individual contributions of our methodology, we conduct systematic ablation studies along two complementary dimensions, with all experiments evaluated on the Biomni-Eval1 benchmark: (1) the effectiveness of the proposed Inner–Outer Dual-Loop architecture, and (2) the impact of staged training strategies, comparing supervised fine-tuning (SFT) and reinforcement learning (RL), on agentic execution quality.
For the Dual-Loop architecture ablation, all architectural variants are evaluated on Biomni-Eval1 using Qwen3-8B-Sci—a model specifically fine-tuned for scientific programming—as a fixed backbone, ensuring that performance differences are solely attributable to architectural design.
For the training strategy ablation, we likewise conduct evaluation on Biomni-Eval1, fixing the backbone to the base Qwen3-8B model and comparing Only SFT against SFT + RL under identical conditions, thereby isolating the effect of reinforcement learning on agentic execution.
@@ 6.1 Ablation on the Dual-Loop Architecture
Table 1: Ablation study of the Dual-Loop Architecture on Biomni-Eval1. Results compare isolated loop configurations against the integrated framework. Best results are highlighted in bold.
@@ 6.2 Ablation on RL-based Staged Training
We further analyze how staged learning enhances the execution quality of the CodeAct module. In this context, staged training refers to an SFT initialization that enforces protocol adherence, followed by RL-based policy optimization that refines decision-making under uncertainty. Using the same Qwen3-8B backbone, we compare the baseline with Only SFT and SFT + RL.
@@
Table 2: Ablation study of RL-based Staged Training on Biomni-Eval1. Best results are highlighted in bold. The baseline takes Qwen-8B as the backbone without post-training.
@@
These results indicate that, in multi-step, multi-tool scientific task settings, the execution layer (the inner-loop in our case) is not merely a passive reasoning endpoint, but a highly learnable and adaptable core component. Through agentic training applied solely to the execution layer, the model achieves substantial improvements in both execution stability and task efficiency, thereby significantly enhancing the reliability of reasoning and task execution for scientific agents in complex scenarios.
This observation further motivates an important research direction: given that complex scientific tasks typically involve long-horizon decision trajectories with extended execution times, it may be a key avenue for advancing the overall capabilities of scientific agents to explore joint optimization mechanisms between the execution layer and the strategic layer within hierarchical agent architectures, to enable cross-level cooperative learning and long-term decision consistency.
@@ 7 Conclusion
Extensive experiments on Biomni-Eval1, ChemBench, and MatSciBench show that S1-NexusAgent achieves state-of-the-art performance across life sciences, chemistry, and materials science, validating both its effectiveness and generalization capability in complex scientific tasks. Case studies further illustrate how the dual-loop architecture and code-centric execution paradigm enable interpretable, adaptive, and resilient scientific workflows in realistic research scenarios.